# Scalar-induced gravitational waves from coherent initial states

### v1

@@ I Introduction
The cosmic microwave background (CMB) remains the most precise probe of the early universe, providing detailed constraints on inflationary physics and cosmological parameters [1, 2, 3]. However, CMB observations are limited to the largest scales — typically beyond megaparsecs — leaving small-scale physics largely unconstrained. In contrast, information encoded in cosmological gravitational waves (GW) remains unconstrained, particularly at much smaller scales, thus providing a window into the small-scale regime complementary to CMB [4, 5, 6, 7, 8, 9, 10, 11, 12]. With the advent of current and proposed GW experiments, the prospect of exploring the early universe at small scales seems increasingly tangible [13, 14, 15, 16, 15, 17, 18, 19, 20, 21].
One of the basic mechanisms for generating a stochastic background of GW is through tensor perturbations sourced by scalar fluctuations at the second order in perturbation theory [22, 23, 24, 25, 26, 11, 12, 27, 28, 29, 30, 31, 32, 7, 33, 34, 35, 36]. These scalar-induced gravitational waves (SIGW) are appealing from a theoretical standpoint, as their production is inevitable and does not rely upon specifics of any model. Furthermore, since SIGW directly couple to the scalar modes, they also encode information about scalar perturbations, possibly at small scales that are otherwise inaccessible to electromagnetic probes. Notably, SIGW potentially observable today can probe modes that exited the horizon during very late stages of inflation, long after those accessible by CMB observation.
@@
In this paper, we investigate the impact of statistical inhomogeneity and anisotropy in primordial scalar perturbations on SIGW. Specifically, we consider a scenario where quantum fluctuations in the inflaton field originate in a coherent state instead of the Bunch-Davies vacuum [55, 56, 57, 54]. For a detailed discussion on the viability and consequence of such a choice of initial quantum states, see [54] and references therein. Quantum fluctuations in a coherent state generally acquire a non-zero one-point function, resulting in a space-dependent mean for the primordial scalar perturbations. The presence of this primordial mean violates the usual assumptions of statistical homogeneity and isotropy of the N-point correlation functions of scalar perturbations. However, the perturbations remain Gaussian. When these primordial fluctuations re-enter the horizon after the end of inflation, they induce tensor modes due to interactions at the second order of perturbation theory. As a result, the tensor perturbations sourced by these scalar modes inherit their statistical features, leading to distinctive signatures in their N-point correlation functions.
The presence of a statistically inhomogeneous primordial scalar mean introduces a one-point function, an induced mean, for the tensor perturbations. Further it modifies the two-point correlation of the induced tensors. Beyond the homogeneous-isotropic power spectrum, this inhomogeneity generates reducible and partially reducible contributions. These contributions enhance the overall spectral density of SIGW. Crucially, the spatial profile of the primordial scalar mean determines the scale dependence of this enhancement. Localized features in the mean can produce a sharp peak in the SIGWs spectral density at specific scales while leaving larger and smaller scales unaffected. Detection of such scale-dependent enhancements in the stochastic GW background could constrain the structure of the primordial mean and therefore, the violation of statistical homogeneity of the primordial perturbations.
@@ II.1 Scalar perturbations
$a_{\bm{k}}\ket{C_{\bm{k}}}$ $=$ ${\rm C}({\bm{k}})|{C_{\bm{k}}}\rangle\,,$ (4)
where $C({\bm{k}})$ is the eigenfunction for a given wavevector ${\bm{k}}$. The coherent state is described by the function $C({\bm{k}})=|C({\bm{k}})|e^{i\theta_{k}}$, where $\theta_{k}$ is the related phase. The number states |$\ket{n_{\bm{k}}}$ are created by $a^{\dagger}_{\bm{k}}$ as $a^{\dagger}_{\bm{k}}\ket{n_{\bm{k}}}=(n_{\bm{k}}+1)\ket{n_{\bm{k}}+1}$ and the lowest state |$\ket{0}$ is the vacuum that is annihilated by $a_{\bm{k}}$ as $a_{\bm{k}}\ket{0}=0$. The explicit expression of |$\ket{C_{\bm{k}}}$ in terms of the number states |$\ket{n_{\bm{k}}}$ is
|$\ket{C_{\bm{k}}}$ $=$ $e^{-\frac{|C({\bm{k}})|^{2}}{2}}\sum_{n=0}^{\infty}C^{n}({\bm{k}})\frac{\ket{n_{\bm{k}}}}{\sqrt{n!}}\,.$ (5)
The operator $D({\bm{k}})$ that generates |$\ket{C_{\bm{k}}}$ from the vacuum |$\ket{0}$ is called the coherent state displacement operator, whose form is [66, 57]
@@
×$\times\Re[\alpha({\bm{k}})]\,\sin({\bm{k}}\cdot{\bm{x}})\,.$
@@ II.2 Tensor perturbations
where ${\bm{e}}_{ij}^{\lambda}({\bm{k}})$ is the polarization tensor for a given polarization index $\lambda$. The index $\lambda$ can be chosen to be +,×+,\,$\times$+ , × states or left and right circular polarization states. The operators $b_{\bm{k}}^{\lambda}$ and $b_{\bm{k}}^{\lambda\,\dagger}$ are the annihilation and creation operators per polarization, obeying the commutation relation
$[b_{{\bm{k}}_{1}}^{\lambda_{1}},b^{\dagger\,\lambda_{2}}_{-{\bm{k}}_{2}}]=\delta^{(3)}({\bm{k}}_{1}+{\bm{k}}_{2})\,\delta^{\lambda_{1}\lambda_{2}}\,.$ (15)
The tensor mode function $g^{\lambda}_{k}(\eta)$ is obtained from the corresponding equation of motion solved over the slow-roll inflationary background. Unlike the scalars, we evolve these tensor perturbations from the Bunch-Davies vacuum |$\ket{0}$ during inflation. While the tensor perturbations can themselves be excited, we restrict our analysis to the case of tensors evolving from the vacuum state. This is done to avoid any confusion between the effects on tensors imprinted by scalars evolving from coherent states and possible features due to tensors evolving from an excited initial state. Therefore the two-point function of $\gamma^{\lambda}_{\bm{k}}$ in the super-Hubble regime is [6, 67]
@@ III Scalar-induced tensor perturbations
The coordinate representation of the polarization factor reveals further symmetries which will be useful for the discussion on chirality of the secondary tensor modes. If we align the vector $\bm{k}$ along the $\hat{\bm{z}}$ axis, $Q^{\lambda}$ in the (+,×)(+$\times$)( + , × ) basis becomes
$Q^{+,\times}(\bm{k},\bm{q})=\frac{1}{\sqrt{2}}q^{2}\sin^{2}\theta\begin{cases}\cos 2\phi&\text{ for }+\\ \sin 2\phi&\text{ for }\times\end{cases},$ (22)
@@
$I(p,q,\eta)=\int_{\eta_{0}}^{\eta}{\rm d}\tilde{\eta}G_{\bm{k}}(\eta,\tilde{\eta})\frac{a(\tilde{\eta})}{a(\eta)}\mathfrak{f}(p,q,\tilde{\eta}),$ (24)
where $G_{\bm{k}}(\eta,\tilde{\eta})$ is the Green function that satisfies
$G^{\prime\prime}_{\bm{k}}(\eta,\tilde{\eta})+\left(k^{2}-\frac{a^{\prime\prime}(\eta)}{a(\eta)}\right)G_{\bm{k}}(\eta,\tilde{\eta})=\delta(\eta-\tilde{\eta}),$ (25)
and $\mathfrak{f}(p,q,\tilde{\eta})$ is sourced by scalar perturbations
@@ III.1 One-point function
×$\quad\times\left[\frac{2\pi^{2}}{q^{3}}\mathcal{P}_{{}_{\mathrm{S}}}(q)\delta^{(3)}(\bm{k})+\Braket{\mathcal{R}_{\bm{q}}}\Braket{\mathcal{R}_{\bm{k-q}}}\right]$ (28)
@@
×$\times\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}-{\bm{q}}|)}{q^{3}|{\bm{k}}-{\bm{q}}|^{3}}}\,\alpha({\bm{q}})\alpha({\bm{k}}-{\bm{q}})\,.$
@@
×$\times\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}-{\bm{q}}|)}{q^{3}|{\bm{k}}-{\bm{q}}|^{3}}}\,\alpha({\bm{q}})\alpha({\bm{k}}-{\bm{q}})\,.$
@@
×$\times\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}-{\bm{q}}|)}{q^{3}|{\bm{k}}-{\bm{q}}|^{3}}}\,\alpha({\bm{q}})\alpha({\bm{k}}-{\bm{q}})\,.$
We can see that $\braket{h^{R,L}_{-{\bm{k}}}}=\braket{h^{R,L}_{\bm{k}}}$ only if $[\alpha({\bm{q}})\alpha({\bm{k}}-{\bm{q}})]^{\ast}=\alpha({\bm{q}})\alpha({\bm{k}}-{\bm{q}})$, i.e. if $\alpha({\bm{k}})$ is a fully real or a fully imaginary function. In case of $\alpha({\bm{k}})$ being a general complex function, we shall have $\braket{h^{\lambda}_{-{\bm{k}}}}\neq\braket{h^{\lambda}_{\bm{k}}}$ indicating violation of parity in tensors, inherited from the sourcing scalars. Thus we have a non-trivial one-point function for the scalar-induced tensor perturbations, that are inhomogeneous, anisotropic and violating parity.
@@ III.2 Two-point function
×$\quad\times I(|\bm{k}_{1}-\bm{q}_{1}|,q_{1},\eta)I(|\bm{k}_{2}-\bm{q}_{2}|,q_{2},\eta)$
×$\quad\times\Braket{\mathcal{R}_{\bm{q}_{1}}\mathcal{R}_{\bm{k}_{1}-\bm{q}_{1}}\mathcal{R}_{\bm{q}_{2}}\mathcal{R}_{\bm{k}_{2}-\bm{q}_{2}}}.$ (32)
@@
$+\Braket{\mathcal{R}_{1}}\Braket{\mathcal{R}_{2}\mathcal{R}_{3}\mathcal{R}_{4}}_{\text{irr}}+\text{2 permutations}$ (33b)
@@
The terms in the first and second line of RHS are zero, since there are no irreducible three- and four-point correlations for Gaussian perturbations. The terms in the third line (33c) are from the four-point function reduced in terms of irreducible two-point functions. Of the three permutations between these terms only two survive. One term contains $\delta^{(3)}({\bm{k}}_{1})\delta^{(3)}({\bm{k}}_{2})$ and so does not contribute to non-zero values of $k_{1}$ or $k_{2}$. These two terms are the statistically homogeneous and isotropic contributions independent of the primordial mean. Next, in the fourth line (33d), there are terms from four-point function reducible to products of one-point and irreducible two-point functions. Once again, of the six terms, two shall contain $\delta^{(3)}({\bm{k}}_{1})\delta^{(3)}({\bm{k}}_{2})$ which do not contribute at non-zero wavenumbers. The last line (33e) is the term that is completely reducible in terms of one-point function.
@@
$\braket{h^{\lambda_{1}}_{{\bm{k}}_{1}}h^{\lambda_{2}}_{{\bm{k}}_{2}}}=$ $f^{\lambda_{1}}(k_{1})\delta^{\lambda_{1}\lambda_{2}}\delta^{(3)}({\bm{k}}_{1}+{\bm{k}}_{2})+g^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2})$
@@
The first term in the RHS above encapsulates contribution from (33c). The presence of $\delta^{(3)}({\bm{k}}_{1}+{\bm{k}}_{2})$ indicates the homogeneous and isotropic nature of this term. The Kronecker delta $\delta^{\lambda_{1}\lambda_{2}}$ implies parity. The explicit form of $f_{\lambda}(k)$ is given by
$f^{\lambda}(k)=32$ $(2\pi^{2})^{2}\int\frac{{\rm d}^{3}{\bm{q}}}{(2\pi)^{3}}|Q^{\lambda}(\bm{k},\bm{q})|^{2}I^{2}(|{\bm{k}}-{\bm{q}}|,q,\eta)$
×$\times\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}-{\bm{q}}|)}{q^{3}|{\bm{k}}-{\bm{q}}|^{3}}\,.$ (35)
The second term of Eq. (34) captures the contribution from (33d). Note that it is neither completely reducible nor irreducible, as it contains terms that mix one-point and two-point functions. The form of $g^{\lambda_{1}\lambda_{2}}(\bm{k}_{1},\bm{k}_{2})$ is given by
$g^{\lambda_{1}\lambda_{2}}(\bm{k}_{1},\bm{k}_{2})=$ $-64(2\pi^{2})^{2}\int\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}Q^{\lambda_{1}}(\bm{k}_{1},\bm{q})Q^{\lambda_{2}}(\bm{k}_{2},\bm{q})$
×$\times I(|\bm{k}_{1}-\bm{q}|,q,\tau)I(|\bm{k}_{2}+\bm{q}|,q,\tau)$
×$\times\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{1}-{\bm{q}}|)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{2}+{\bm{q}}|)}{|{\bm{k}}_{1}-{\bm{q}}|^{3}|{\bm{k}}_{2}+{\bm{q}}|^{3}}}\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)}{q^{3}}$
×$\times\alpha({\bm{k}}_{1}-{\bm{q}})\alpha({\bm{k}}_{2}+{\bm{q}})\,.$ (36)
@@
×$\times Q^{\lambda_{1}}({\bm{k}}_{1},{\bm{q}}_{1})\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q_{1})\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{1}-{\bm{q}}_{1}|)}{q_{1}^{3}|{\bm{k}}_{1}-{\bm{q}}_{1}|^{3}}}$
×$\times\alpha({\bm{q}}_{1})\alpha({\bm{k}}_{1}-{\bm{q}}_{1})\Bigg{]}$
×$\times\Bigg{[}\int\frac{{\rm d}^{3}{\bm{q}}_{2}}{(2\pi)^{3/2}}Q^{\lambda_{2}}({\bm{k}}_{2},{\bm{q}}_{2})I(|{\bm{k}}_{2}-{\bm{q}}_{2}|,q_{2},\eta)$
×$\times\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q_{2})\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{2}-{\bm{q}}_{2}|)}{q^{3}|{\bm{k}}_{2}-{\bm{q}}_{2}|^{3}}}\,\alpha({\bm{q}}_{2})\alpha({\bm{k}}_{2}-{\bm{q}}_{2})\Bigg{]}\,.$

### v2

@@ I Introduction
The cosmic microwave background (CMB) remains the most precise probe of the early universe, providing detailed constraints on inflationary physics and cosmological parameters [1, 2, 3]. However, CMB observations are limited to the largest scales — typically beyond megaparsecs — leaving small-scale physics largely unconstrained. In contrast, information encoded in cosmological gravitational waves (GW) remains unconstrained, particularly at much smaller scales, thus providing a window into the small-scale regime complementary to CMB [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. With the advent of current and proposed GW experiments, the prospect of exploring the early universe at small scales seems increasingly tangible [15, 16, 17, 18, 17, 19, 20, 21, 22, 23].
One of the basic mechanisms for generating a stochastic background of GW is through tensor perturbations sourced by scalar fluctuations at the second order in perturbation theory [24, 25, 26, 27, 28, 11, 12, 29, 30, 31, 32, 33, 34, 7, 35, 36, 37, 38, 39, 40]. These scalar-induced gravitational waves (SIGW) are appealing from a theoretical standpoint, as their production is inevitable and does not rely upon specifics of any model. Furthermore, since SIGW directly couple to the scalar modes, they also encode information about scalar perturbations, possibly at small scales that are otherwise inaccessible to electromagnetic probes. Notably, SIGW potentially observable today can probe modes that exited the horizon during very late stages of inflation, long after those accessible by CMB observation.
@@
In this paper, we investigate the impact of statistical inhomogeneity and anisotropy in primordial scalar perturbations on SIGW. Specifically, we consider a scenario where quantum fluctuations in the inflaton field originate in a coherent state instead of the Bunch-Davies vacuum [59, 60, 61, 62, 58]. For a detailed discussion on the viability and consequence of such a choice of initial quantum states, see [58] and references therein. Quantum fluctuations in a coherent state generally acquire a non-zero one-point function, resulting in a space-dependent mean for the primordial scalar perturbations. The presence of this primordial mean violates the usual assumptions of statistical homogeneity and isotropy of the N-point correlation functions of scalar perturbations. However, the perturbations remain Gaussian. When these primordial fluctuations re-enter the horizon after the end of inflation, they induce tensor modes due to interactions at the second order of perturbation theory. As a result, the tensor perturbations sourced by these scalar modes inherit their statistical features, leading to distinctive signatures in their N-point correlation functions.
The presence of a statistically inhomogeneous primordial scalar mean introduces a one-point function, an induced mean, for the tensor perturbations. Further it modifies the two-point correlation of the induced tensors. Beyond the homogeneous-isotropic power spectrum, this inhomogeneity generates reducible and partially reducible contributions. These contributions enhance the overall spectral density of SIGW. Crucially, the spatial profile of the primordial scalar mean determines the scale dependence of this enhancement. Localized features in the mean can produce a sharp peak in the SIGW’s spectral density at specific scales while leaving larger and smaller scales unaffected. Detection of such scale-dependent enhancements in the stochastic GW background could constrain the structure of the primordial mean and therefore, the violation of statistical homogeneity of the primordial perturbations.
@@ II.1 Scalar perturbations
[…]
@@
$\times\Re[\alpha({\bm{k}})]\,\sin({\bm{k}}\cdot{\bm{x}})\,.$
@@ II.2 Tensor perturbations
where ${\bm{e}}_{ij}^{\lambda}({\bm{k}})$ is the polarization tensor for a given polarization index $\lambda$. The index $\lambda$ can be chosen to be $\,\times$ states or left and right circular polarization states. The operators $b_{\bm{k}}^{\lambda}$ and $b_{\bm{k}}^{\lambda\,\dagger}$ are the annihilation and creation operators per polarization, obeying the commutation relation
$b_{{\bm{k}}_{1}}^{\lambda_{1}},b^{\dagger\,\lambda_{2}}_{-{\bm{k}}_{2}}]=\delta^{(3)}({\bm{k}}_{1}+{\bm{k}}_{2})\,\delta^{\lambda_{1}\lambda_{2}}\,.$ (15)
[…]
@@ III Scalar-induced tensor perturbations
The coordinate representation of the polarization factor reveals further symmetries which will be useful for the discussion on chirality of the secondary tensor modes. If we align the vector $\bm{k}$ along the $\hat{\bm{z}}$ axis, $Q^{\lambda}$ in the $(+,\times)$ basis becomes
[…]
@@
[…]
@@ III.1 One-point function
$\quad\times\left[\frac{2\pi^{2}}{q^{3}}\mathcal{P}_{{}_{\mathrm{S}}}(q)\delta^{(3)}(\bm{k})+\Braket{\mathcal{R}_{\bm{q}}}\Braket{\mathcal{R}_{\bm{k-q}}}\right]$ (28)
@@
$\times\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}-{\bm{q}}|)}{q^{3}|{\bm{k}}-{\bm{q}}|^{3}}}\,\alpha({\bm{q}})\alpha({\bm{k}}-{\bm{q}})\,.$
@@
$\times\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}-{\bm{q}}|)}{q^{3}|{\bm{k}}-{\bm{q}}|^{3}}}\,\alpha({\bm{q}})\alpha({\bm{k}}-{\bm{q}})\,.$
@@
$\times\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}-{\bm{q}}|)}{q^{3}|{\bm{k}}-{\bm{q}}|^{3}}}\,\alpha({\bm{q}})\alpha({\bm{k}}-{\bm{q}})\,.$
We can see that $\braket{h^{R,L}_{-{\bm{k}}}}=\braket{h^{R,L}_{\bm{k}}}$ only if $\alpha({\bm{q}})\alpha({\bm{k}}-{\bm{q}})]^{\ast}=\alpha({\bm{q}})\alpha({\bm{k}}-{\bm{q}})$, i.e. if $\alpha({\bm{k}})$ is a fully real or a fully imaginary function. In case of $\alpha({\bm{k}})$ being a general complex function, we shall have $\braket{h^{\lambda}_{-{\bm{k}}}}\neq\braket{h^{\lambda}_{\bm{k}}}$ indicating violation of parity in tensors, inherited from the sourcing scalars. Thus we have a non-trivial one-point function for the scalar-induced tensor perturbations, that are inhomogeneous, anisotropic and violating parity.
@@ III.2 Two-point function
$\quad\times I(|\bm{k}_{1}-\bm{q}_{1}|,q_{1},\eta)I(|\bm{k}_{2}-\bm{q}_{2}|,q_{2},\eta)$
$\quad\times\Braket{\mathcal{R}_{\bm{q}_{1}}\mathcal{R}_{\bm{k}_{1}-\bm{q}_{1}}\mathcal{R}_{\bm{q}_{2}}\mathcal{R}_{\bm{k}_{2}-\bm{q}_{2}}}.$ (32)
@@
$+\Braket{\mathcal{R}_{1}}\Braket{\mathcal{R}_{2}\mathcal{R}_{3}\mathcal{R}_{4}}_{\text{irr}}+\text{3 permutations}$ (33b)
@@
The terms in the first and second line of RHS are zero, since there are no irreducible three- and four-point correlations for Gaussian perturbations. The terms in the third line (33c) are from the four-point function reduced in terms of irreducible two-point functions. Of the three permutations between these terms only two survive. One term contains $\delta^{(3)}({\bm{k}}_{1})\delta^{(3)}({\bm{k}}_{2})$ and so does not contribute to non-zero values of $k_{1}$ or $k_{2}$. These two terms are the statistically homogeneous and isotropic contributions independent of the primordial mean. Next, in the fourth line (33d), there are terms from four-point function reducible to products of one-point and irreducible two-point functions. Once again, of the six terms, two shall contain $\delta^{(3)}({\bm{k}}_{1})$ and $\delta^{(3)}({\bm{k}}_{2})$ which do not contribute at non-zero wavenumbers. The last line (33e) is the term that is completely reducible in terms of one-point function.
@@
$\braket{h^{\lambda_{1}}_{{\bm{k}}_{1}}h^{\lambda_{2}}_{{\bm{k}}_{2}}}=$ $f^{\lambda_{1}}(k_{1},\eta)\delta^{\lambda_{1}\lambda_{2}}\delta^{(3)}({\bm{k}}_{1}+{\bm{k}}_{2})+g^{\lambda_{1}\lambda_{2}}({\bm{k}}_{1},{\bm{k}}_{2},\eta)$
@@
The first term in the RHS above encapsulates contribution from (33c). The presence of $\delta^{(3)}({\bm{k}}_{1}+{\bm{k}}_{2})$ indicates the homogeneous and isotropic nature of this term. The Kronecker delta $\delta^{\lambda_{1}\lambda_{2}}$ implies parity. The explicit form of $f_{\lambda}(k{,\eta})$ is given by
$f^{\lambda}(k{,\eta})=32$ $(2\pi^{2})^{2}\int\frac{{\rm d}^{3}{\bm{q}}}{(2\pi)^{3}}|Q^{\lambda}(\bm{k},\bm{q})|^{2}I^{2}(|{\bm{k}}-{\bm{q}}|,q,\eta)$
$\times\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}-{\bm{q}}|)}{q^{3}|{\bm{k}}-{\bm{q}}|^{3}}\,.$ (35)
The second term of Eq. (34) captures the contribution from (33d). Note that it is neither completely reducible nor irreducible, as it contains terms that mix one-point and two-point functions. The form of $g^{\lambda_{1}\lambda_{2}}(\bm{k}_{1},\bm{k}_{2}{,\eta})$ is given by
$g^{\lambda_{1}\lambda_{2}}(\bm{k}_{1},\bm{k}_{2}{,\eta})=$ $-64(2\pi^{2})^{2}\int\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}Q^{\lambda_{1}}(\bm{k}_{1},\bm{q})Q^{\lambda_{2}}(\bm{k}_{2},\bm{q})$
$\times I(|\bm{k}_{1}-\bm{q}|,q,\eta)I(|\bm{k}_{2}+\bm{q}|,q,\eta)$
$\times\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{1}-{\bm{q}}|)\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{2}+{\bm{q}}|)}{|{\bm{k}}_{1}-{\bm{q}}|^{3}|{\bm{k}}_{2}+{\bm{q}}|^{3}}}\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q)}{q^{3}}$
$\times\alpha({\bm{k}}_{1}-{\bm{q}})\alpha({\bm{k}}_{2}+{\bm{q}})\,.$ (36)
@@
$\times Q^{\lambda_{1}}({\bm{k}}_{1},{\bm{q}}_{1})\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q_{1})\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{1}-{\bm{q}}_{1}|)}{q_{1}^{3}|{\bm{k}}_{1}-{\bm{q}}_{1}|^{3}}}$
$\times\alpha({\bm{q}}_{1})\alpha({\bm{k}}_{1}-{\bm{q}}_{1})\Bigg]$
$\times\Bigg[\int\frac{{\rm d}^{3}{\bm{q}}_{2}}{(2\pi)^{3/2}}Q^{\lambda_{2}}({\bm{k}}_{2},{\bm{q}}_{2})I(|{\bm{k}}_{2}-{\bm{q}}_{2}|,q_{2},\eta)$
$\times\sqrt{\frac{\mathcal{P}_{{}_{\mathrm{S}}}(q_{2})\mathcal{P}_{{}_{\mathrm{S}}}(|{\bm{k}}_{2}-{\bm{q}}_{2}|)}{q^{3}|{\bm{k}}_{2}-{\bm{q}}_{2}|^{3}}}\,\alpha({\bm{q}}_{2})\alpha({\bm{k}}_{2}-{\bm{q}}_{2})\Bigg]\,.$